OpenStudy (anonymous):

use f(x)=3x-4 and g(x)=2x^2 to evaluate each of the following (a)(f+g)(3) (b)(f-g)(-2) (c)(fg)(7) (d)(f/g)(0)

OpenStudy (anonymous):

(a)(f+g)(3)=(3x-4)+(2x^2) =(3(3)-4)+(2(3)^2) =(9-4)+(2*9) =5+18 = 23

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

(b)(f-g)(-2)=(3x-4)-(2x^2) =(3(-2)-4)-(2(-2)^2) =-10-8 =-18

OpenStudy (anonymous):

thank you so much I am so lost in this that I did not know there were two answers.

OpenStudy (anonymous):

(c)(fg)(7)=(3(7)-4)*(2(7)^2) =(21-4)(2*49) = 17*98 =1.666

OpenStudy (anonymous):

(d)(f/g)(0)=\[\frac{3x-4}{2x^2}=\frac{3(0)-4}{2(0)^2}=\frac{0-4}{0}=0\]

OpenStudy (anonymous):

if u like my answer ,click good answer,ty